Help I'm completely lost on this!
Determine if each relation is Linear or Not Linear.
x y
-2 -14
3 11
5 21

Answer by stanbon(75887)

Determine if each relation is Linear or Not Linear.
x y
-2 -14
3 11
5 21
----------
Check the slope between successive points:
[11--14]/[3--2] = 25/5 = 5
[21-11]/[5-3] = 10/2 = 5
[21--14]/[5--2] = 35/7 = 5
---------------
Slopes are the same; the relation is linear function.
======================
Cheers,
Stan H.